OpenStudy (anonymous):

how to simplify (x-3)(x^2-4x-7) step by step

OpenStudy (anonymous):

first distribute the x-3 to every term in the other set of parenthesis

hero (hero):

(x - 3)(x^2 - 4x - 7) = x(x^2 - 4x - 7) - 3(x^2 - 4x - 7) = x^3 - 4x^2 - 7x - 3x^2 + 12x + 21 Finish it

OpenStudy (anonymous):

Thank you so much.!:)

OpenStudy (anonymous):

hey Hero would you mind helping me with my question?

hero (hero):

You can also do it this way: (x - 3)(x^2 - 4x - 7) = x^2(x - 3) - 4x(x - 3) - 7(x - 3) = x^3 - 3x^2 - 4x^2 + 12x - 7x - 21

hero (hero):

Both approaches are applications of the distributive property